Allen bought 20 stamps at the post office in 37¢ and 20¢ denominations. If the total cost of the stamps was $7.06, how many 37¢ stamps did Allen buy?
Answer by venugopalramana(3286)

Allen bought 20 stamps at the post office in 37¢(X SAY) and 20¢ (Y SAY)denominations. If the total cost of the stamps was $7.06, how many 37¢ stamps
did Allen buy?
X+Y=20.....Y=20-X......I
VALUE =37X+20Y=706..........II
37X+20(20-X)=706
37X+400-20X=706
17X=706-400=306
X=306/17=18